OpenStudy (anonymous):

solve for B: V = 1/3BH

OpenStudy (anonymous):

3bvh=1 bvh=1/3 bh=1/3v b=1/3vh

OpenStudy (anonymous):

thanks

OpenStudy (ajprincess):

 1.Multiply both side by 3. 2.Divide by H both sides

OpenStudy (anonymous):

I'm pretty sure you mean \[V=\frac{1}{3}bh\] Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):